OpenStudy (anonymous):

write the standard form of the equation of the circle with endpoints of a diameter with the points (1,8) and (-3,11)

OpenStudy (anonymous):

center is the midpoint of the line joining the two points. midpoint is \[(\frac{1-3}{2},\frac{8+11}{2})\] \[(-1,\frac{19}{2})\]

OpenStudy (anonymous):

lalaly will complete it

OpenStudy (lalaly):

find the mid point x= (1+-3)/2=-1 y=(8+11)/2=19/2 find the distance squareroot of (-3-8)^2+(11-8)^2 = sqrt of 121+9= square root of 130 so radius is (squareroot of 130)/2 equation is (x+1)^2+(y-19/2)^2=130/4

OpenStudy (anonymous):

thank you!

OpenStudy (lalaly):

;);)